In a given amount of time Van can produce either 50 pounds of vegetables or 10 pounds of chicken.

If Van wants to produce 10 pounds of chicken then he must give up ____pounds of vegetables.

If Van wants to produce 1 pound of chicken then he must give up ____ pounds of vegetables.

If Van wants to produce 2 pounds of chicken then he must give up ____ pounds of vegetables.
Answer #1

Considering the opportunity cost o producing chicken as : C/V = 10/50 = 5

Thus for each pound of chicken Van must give up 5 pounds of Vegetables

f Van wants to produce 10 pounds of chicken then he must give up 50 pounds of vegetables.

If Van wants to produce 1 pound of chicken then he must give up 5 pounds of vegetables.

If Van wants to produce 2 pounds of chicken then he must give up 10 pounds of vegetables.